# Strategic Classification with Non-Linear Classifiers

### v1

@@ C.4 Non-bijectiveness of Classifier Mapping
In Sec. 5.3, we show that the mapping $h\mapsto h_{\Delta}$ is not surjective by showing that there are many potential $h_{\Delta}$ with no associated $h$. Here, we will show that $h\mapsto h_{\Delta}$ is also not injective since many classifiers map to the same effective classifier. Furthermore, we show that there are actually infinite $h$ that map to nearly all possible $h_{\Delta}$. Though there are many more interesting examples, the simplest case of non-injectiveness can be seen by considering an $h$ that includes a continuous positive region $C$, where $|C|=\infty$. Define the classifier $h^{x}$ as the classifier that is identical to $h$ except that $h^{x}$ classifies $x\in C$ as negative instead of positive. Because point $x$ will achieve a positive classification through strategic movement, $h^{x}_{\Delta}$ is unaffected by $x$ and is the same as $h_{\Delta}$. Moreover, $h^{x}_{\Delta}=h_{\Delta}$ is true of any $x\in C$, so as long as any $h$ that maps to $h_{\Delta}$ contains a positive region with infinite points, then $h_{\Delta}$ will have infinite $h$ that all map to it. Because positive points in $h$ map to positive radius-$\alpha$ balls in $h_{\Delta}$, any $h_{\Delta}$ with a positive region $C$ such that $\exists x:\,\,B_{\alpha}(x)\subset C$ will have infinite $h$ that all map to it. Note that since all positive regions in $h_{\Delta}$ have a minimum size requirement $\exists x:\,\,B_{\alpha}(x)\subseteq C$, the aforementioned requirement covers all but the $h_{\Delta}$ with absolute minimum-size positive regions.
C.5 Classes with $H=H_{\Delta}$
@@
Though $H\not=H_{\Delta}$ for most classes $H$, there are a few rather simple $H$ other than the class of linear functions where $H=H_{\Delta}$. For example, under any $\ell_{p}$ norm cost, an $\ell_{p}$ ball with a positive inside will grow by a radius of $\alpha$, while an $\ell_{p}$ ball with a negative inside will shrink by a radius of $\alpha$ (at least until the radius hits 0). Therefore, many model classes of the subsets of the $\ell_{p}$ balls – including the entire class of $\ell_{p}$ balls , the class of all negative-inside balls, and the class of balls with a radius that is a multiple of $\alpha$ – all have $H=H_{\Delta}$ for any fixed $\alpha$.
@@
Additionally, since each line segment and $\ell_{p}$ arc maps to another line segment or $\ell_{p}$ arc, many peicewise-linear or peicewise $\ell_{p}$ arc classes will also have $H=H_{\Delta}$. For example, any concave linear intersection will be mapped to a shifted concave linear intersection, so the class of concave linear intersection satisfies $H=H_{\Delta}$.
C.6 Classes with $\mathrm{VC}(H)>\mathrm{VC}(H_{\Delta})>0$
@@
Figure 11: Example of VC decreasing from 2 to 1. Though the black points can be shattered by $H$, none of the $h_{\Delta}$ overlap so $\mathrm{VC}(H_{\Delta})=1$
@@
There are a couple of mechanisms that can lead to a VC decrease. In the first mechanism, multiple different $h$ have the same $h_{\Delta}$ leading to a smaller and less expressive $H_{\Delta}$. Two such cases are shown in Appendix B.8 to prove that the VC may decrease even from $\infty$ to 0. In a similar fashion, we can build an $H$ where only some (instead of all) of the $h\in H$ have the same $h_{\Delta}$, which would decrease the VC dimension to a non-zero effective VC dimension.
@@
In the second mechanism, the VC can change due to effective classifiers gaining or losing overlap from strategic movement. As the dual to Fig. 5 (Left) where the VC increases from 1 to 2 due to this mechanism, Fig. 11 presents an example where the VC decreases from 2 to 1. Here, H includes 4 overlapping classifiers with $\mathrm{VC}(H)=2$, yet under strategic behavior, the effective classifiers cease to overlap decreasing the shattering capacity to 1.
C.7 Example of $H_{\Delta}$ with $\mathrm{VC}(H_{\Delta})=\Theta(d\cdot(\mathrm{VC}(H))$
@@
For simplification of notation, we build on the case shown in Fig. 5 (Left) using classifiers $h$ that classify some ball as positive and all other points as negative. Note that a similar proof can be built based on a class $H$ of point classifiers [7], but we have decided to include the following construction since classifiers that classify only a single point as positive are quite unrealistic. Let set $S=\mathbb{Z}_{2}^{d}-\{0^{d}\}+\{-\delta^{d}\}$ be the set of $\mathbb{Z}_{2}^{d}$ lattice points with the point $\{0^{d}\}$ replaced by the point $\{-\delta^{d}\}$. Consider the case where $\alpha=0.75$ and $H=\{B_{0.25}(x):x\in S\}$ is the class of radius-0.25 balls around the points in $S$. Because the radius of positive balls expand by $\alpha$ under strategic movement, we get that $H_{\Delta}=\{B_{1}(x):x\in S\}$. Note that because none of the balls in $H$ overlap, $\mathrm{VC}(H)=1$. On the other hand, the balls in $H_{\Delta}$ do overlap allowing $H_{\Delta}$ to shatter the set of $d$ one-hot points $E=\{e_{i}:i\in[1...d]\}$, where $e_{i}$ denotes the point with a 1 in the ith coordinate and 0’s elsewhere. Note that for each $x\in S-\{-\delta^{d}\}$, the ball $B_{1}(x)\in H_{\Delta}$ classifies $E$ according to the sign pattern $x$. Additionally, the ball $B_{1}(-\delta^{d})$ does not classify any point in $E$ as positive, so $H_{\Delta}$ covers all $2^{d}$ sign patterns of $E$. As such, $\mathrm{VC}(H_{\Delta})\geq d$. However, since $H_{\Delta}\subset B=\{B_{1}(x):x\in R^{d}\}$, we get that $\mathrm{VC}(H_{\Delta})\leq\mathrm{VC}(B)=d+1$ [15] giving $\mathrm{VC}(H_{\Delta})=\Theta(d\cdot(VC(H))$. ∎
C.8 Example of Strategic Accuracy Exceeding Standard Accuracy
@@
In the unrealizable setting, the the maximum strategic accuracy can exceed the maximum standard accuracy when strategic movement helps correct the incorrect classification of positive points. For example, consider a basic setup where $\alpha=1$, $X=\{(-1,0),(1,0),(3,0)\}$, $Y=\{0,0,1\}$, and $H=\{h_{1}(x)=sign(x_{1}+x_{2}),h_{2}(x)=sign(x_{1}-x_{2})\}$. In the standard case, both $h_{1}$ and $h_{2}$ correctly classify the first and third points, but not the second point. However, in the strategic setting, the second point is able to obtain a positive classification under either classifier, so the maximum strategic accuracy is 1. ∎
@@ D.1 Experiment #1: Expressivity
In Sec. 7.1, we experimentally demonstrate how the strategic setting affects classifier expressivity. In this experiment, we first randomly sample a degree-$k$ polynomial $h$, then determine the set of points $S$ that make up the decision boundary of $h_{\Delta}$ (i.e. are directly adjacent to it), and finally find the best approximate polynomial fit of $S$. Because user features are often naturally bounded, we assume that $x\in{\cal{X}}=\mathbb{R}^{2}$ and that all $x$ are bounded by $|x|_{\infty}\leq 100$. We report average values and standard error confidence intervals of 100 instances of the setup for each $\alpha\in\{2,4,8,16,24,32,40\}$ and $k\in[1,10]$. The instances were divided among 100 CPUs to speed up computation.
To randomly sample polynomials of degree $k$, we leverage the fact that the number of points needed to uniquely determine a degree-$k$ polynomial over $\mathbb{R}^{2}$ is ($k+2\choose 2$. Therefore, for each instance, we randomly choose ($k+2\choose 2$ points and labels, and use an SVM polynomial fitter to find the best fitting degree-$k$ polynomial fit to these points. To assure that the generated polynomial is indeed a full-degree polynomial, we construct a grid $G$ of points over $[-100,100]^{2}$, label each point $g\in G$ using $h$ to get labels $Y$, and verify that the SVM best-fitting degree-$k$ polynomial achieves near perfect accuracy on ($G,Y$), while the best $(k-1)$-degree polynomial achieves low accuracy on ($G,Y$).
@@ D.2 Experiment #2: Approximation
In Sec. 7.2, we experimentally demonstrate the effects non-universality on the maximum strategic accuracy. This experiment consists of (i) sampling synthetic data and (ii) calculating the maximum linear accuracy and strategic accuracy on each dataset instance. We report average values and standard error confidence intervals of 20 instances of the setup for each $\alpha\in\{0.5,1,2\}$ and $\mu\in[0,5]$. The instances were divided among 100 CPUs to speed up computation.
@@
To calculate the maximum linear accuracy in each instance, we first note that there exists an optimal linear classifier that runs through (or infinitesimally close to for negative points) exactly two dataset points. This is because for any optimal classifier $h$ that runs through exactly one dataset point $x\in{\cal{X}}$, $h$ may be rotated until it runs through another $x^{\prime}\in{\cal{X}}$ without changing any of the labels. For any optimal classifier $h$ that runs through no $x\in{\cal{X}}$, $h$ can be shifted until it runs through one (or two collinear) $x\in{\cal{X}}$ and then rotated appropriately. Because all data points are random, we assume that no three points are collinear. Therefore, the maximum linear accuracy on ${\cal{X}}$ can be calculated by taking the maximum accuracy over the ($n\choose 2$ classifiers that run through each pair of $x,x^{\prime}\in{\cal{X}}$  [^8] Because points on the decision boundary are labeled as positive, we manually include that negative points used to define the linear classifier are accurately labeled to reflect the fact that the optimal classifier does not run through the point, but infinitely close to it..
When calculating the maximum strategic accuracy, we note that the minimum size of a positive region in $h_{\Delta}$ is the ball $B_{\alpha}$, so any strategic classifier can be represented by the intersection of multiple $B_{\alpha}$. We therefore calculate all possible subsets $s\subseteq{\cal{X}}$ that any $B_{\alpha}$ classifies as positive and find the intersection $\bigcap_{s\in S^{\prime}\subseteq S}\,s$ with the highest accuracy. To find all $s\in S$, we iterate through the circles centered at each point on a fine-grained grid over the range of the dataset (padded by $\alpha$ on all sides) and check which points $x\in{\cal{X}}$ each circle includes. To find the best intersection of these circles, we iterate through each of the subsets of $Z\subseteq{\cal{X}}_{0}$  [^9] Subsets are traversed in size order with early stopping once the size of the subset is large enough that the accuracy will be worse than the current best accuracy even if all positive points are correctly classified. and then $s\in S$ to find the maximum number of positive points that can be classified as positive if we allow ourselves to mistakenly classify the negative points in $Z$.

### v2

@@ C.5 Non-bijectiveness of Classifier Mapping
In Sec. 4.4, we show that the mapping $h\mapsto h_{\Delta}$ is not surjective by showing that there are many potential $h_{\Delta}$ with no associated $h$. Here, we will show that $h\mapsto h_{\Delta}$ is also not injective since many classifiers map to the same effective classifier. Furthermore, we show that there are actually infinite $h$ that map to nearly all possible $h_{\Delta}$. Though there are many more interesting examples, the simplest case of non-injectiveness can be seen by considering an $h$ that includes a continuous positive region $C$ with an infinite number of points. Define the classifier $h^{x}$ as the classifier that is identical to $h$ except that $h^{x}$ classifies ${x}\in C$ as negative instead of positive. Because the point x{x} will achieve a positive classification through strategic movement, $h^{x}_{\Delta}$ is unaffected by x{x} and is the same as $h_{\Delta}$. Moreover, $h^{x}_{\Delta}=h_{\Delta}$ is true of any ${x}\in C$, so as long as any $h$ that maps to $h_{\Delta}$ contains a positive region with infinite points, then $h_{\Delta}$ will have infinite $h$ that all map to it. Because positive points in $h$ map to positive radius-$\alpha$ balls in $h_{\Delta}$, any $h_{\Delta}$ with a positive region $C$ such that $\exists x:\,\,B_{\alpha}(x)\subset C$ will have infinite $h$ that all map to it. Note that since all positive regions in $h_{\Delta}$ have a minimum size requirement $\exists x:\,\,B_{\alpha}(x)\subseteq C$, the aforementioned requirement covers all but the $h_{\Delta}$ with absolute minimum-size positive regions.
C.6 Classes with $H=H_{\Delta}$
@@
Though $H\not=H_{\Delta}$ for most classes $H$, there are a few rather simple $H$ other than the class of linear functions where $H=H_{\Delta}$. For example, under any $\ell_{p}$ norm cost, an $\ell_{p}$ ball with a positive inside will grow by a radius of $\alpha$, while an $\ell_{p}$ ball with a negative inside will shrink by a radius of $\alpha$ (at least until the radius hits 0). Therefore, many model classes of the subsets of the $\ell_{p}$ balls — including the class of all negative-inside balls and the class of balls with a radius that is a multiple of $\alpha$ — all have $H=H_{\Delta}$ for any fixed $\alpha$.
@@
Additionally, since each line segment and $\ell_{p}$ arc maps to another line segment or $\ell_{p}$ arc, many piecewise-linear or piecewise $\ell_{p}$ arc classes will also have $H=H_{\Delta}$. For example, any concave linear intersection will be mapped to a shifted concave linear intersection, so the class of concave linear intersections satisfies $H=H_{\Delta}$.
C.7 Classes with $\mathrm{VC}(H)>\mathrm{VC}(H_{\Delta})>0$
@@
Figure 11: Example of VC decreasing from 2 to 1. Each color represents a different classifier (drawn together to illustrate shattering capacity), and the dashed curves represent the corresponding effective classifiers. Though the black points can be shattered by $H$, none of the $h_{\Delta}$ overlap so $\mathrm{VC}(H_{\Delta})=1$.
@@
There are a couple of mechanisms that can lead to a decrease in VC dimension. In the first mechanism, multiple different $h$ have the same $h_{\Delta}$ leading to a smaller and less expressive $H_{\Delta}$. Two such cases are shown in Appendix B.8 to prove that the VC may decrease even from $\infty$ to 0. In a similar fashion, we can build an $H$ where only some (instead of all) of the $h\in H$ have the same $h_{\Delta}$, which would decrease the VC dimension to a non-zero effective VC dimension.
@@
In the second mechanism, the VC can change due to effective classifiers gaining or losing overlap from strategic movement. As the dual to Fig. 4 (Left) where the VC increases from 1 to 2 due to this mechanism, Fig. 11 presents an example where the VC decreases from 2 to 1. Here, $H$ includes four overlapping classifiers with $\mathrm{VC}(H)=2$. Yet, under strategic behavior, the effective classifiers cease to overlap, decreasing the shattering capacity to 1.
C.8 Example of $H_{\Delta}$ with $\mathrm{VC}(H_{\Delta})=\Theta(d\cdot(\mathrm{VC}(H))$
@@
For simplification of notation, we build on the case shown in Fig. 4 (Left) using classifiers $h$ that classify some ball as positive and all other points as negative. Note that a similar proof can be built based on a class $H$ of point classifiers [7], but we have decided to include the following construction since classifiers that classify only a single point as positive are quite unrealistic. Let set $S=\mathbb{Z}_{2}^{d}-\{0^{d}\}+\{-\delta^{d}\}$ be the set of $\mathbb{Z}_{2}^{d}$ lattice points with the point $\{0^{d}\}$ replaced by the point $\{-\delta^{d}\}$. Consider the case where $\alpha=0.75$ and $H=\{B_{0.25}(x):x\in S\}$ is the class of radius-0.25 balls around the points in $S$. Because the radius of positive balls expand by $\alpha$ under strategic movement, we get that $H_{\Delta}=\{B_{1}(x):x\in S\}$. Note that because none of the balls in $H$ overlap, $\mathrm{VC}(H)=1$. On the other hand, the balls in $H_{\Delta}$ do overlap, allowing $H_{\Delta}$ to shatter the set of $d$ one-hot points $E=\{e_{i}:i\in[1...d]\}$, where $e_{i}$ denotes the point with a 1 in the ith coordinate and 0’s elsewhere. Note that for each $x\in S-\{-\delta^{d}\}$, the ball $B_{1}(x)\in H_{\Delta}$ classifies $E$ according to the sign pattern $x$. Additionally, the ball $B_{1}(-\delta^{d})$ does not classify any point in $E$ as positive, so $H_{\Delta}$ covers all $2^{d}$ sign patterns of $E$. As such, $\mathrm{VC}(H_{\Delta})\geq d$. However, since $H_{\Delta}\subset B=\{B_{1}(x):x\in R^{d}\}$, we get that $\mathrm{VC}(H_{\Delta})\leq\mathrm{VC}(B)=d+1$ [15] giving $\mathrm{VC}(H_{\Delta})=\Theta(d\cdot(VC(H))$. ∎
C.9 Example of Strategic Accuracy Exceeding Standard Accuracy
@@
In the unrealizable setting, the maximum strategic accuracy can exceed the maximum standard accuracy when strategic movement helps correct the incorrect classification of positive points. For example, consider a basic setup where $\alpha=1$, $X=\{(-1,0),(2,0),(-3,0)\}$, $Y=\{1,1,-1\}$, and $H=\{h_{1}(x)=sign(x_{1}+x_{2}),h_{2}(x)=sign(x_{1}-x_{2}-1)\}$. In the standard case, both $h_{1}$ and $h_{2}$ correctly classify the second and third points, but not the first point, so the maximum accuracy is $\frac{2}{3}$. However, in the strategic setting, the first point is able to obtain a positive classification under $h_{1}$ by moving to $(-\frac{1}{3},\frac{2}{3})$ (and the third point still cannot get a positive prediction), so the maximum strategic accuracy is 1. ∎
C.10 Practical Takeaways
Given the challenge of optimization in strategic batch learning (even in simpler settings), our paper aims to first establish an understanding of the challenges inherent to non-linear strategic learning (in particular in relation to the linear case), with the hope that our results and conclusions can help guide the future design of learning algorithms, as well as set expectations for what is achievable and what is not. The difficulty of designing a principled algorithm for the general non-linear case is underscored by the fact that existing works rely on strong assumptions to enable optimization. For example, the original paper of Hardt et al. [16] provides an algorithm, but requires assumptions on the cost function that reduce the problem to a one-dimensional learning task, and their algorithm is essentially a line search. Levanon and Rosenfeld [25] differentiate through $\Delta$, but this relies strongly on this operation being a linear projection, which applies only to linear classifiers. Neither of the above naturally extend to the non-linear case.
Towards the goal of practical learning in the general strategic setting, we provide here some examples of how our results can be used as practical takeaways for effective learning:
  1. 1.
    Learning via inversion: One implication of our function-level analysis is that for any classifier $h$ with strategic accuracy $\mathrm{acc}_{strat}(h)$, the classifier $h^{\prime}$ such that $h\mapsto h_{\Delta}=h^{\prime}$ has regular (i.e., non-strategic) accuracy $\mathrm{acc}(h^{\prime})=\mathrm{acc}_{strat}(h)$. Thus, one approach to optimizing strategic accuracy is via reduction: (i) train $h^{\prime}\in H^{\prime}$ to maximize non-strategic accuracy using any conventional approach for some choice of $H^{\prime}$, and then (ii) apply the inverse function mapping (which may not be 1-to-1) to obtain a non-strategic classifier $h$, whose effective decision boundary is that of $h^{\prime}$. This approach requires the ability to solve the ‘inverse’ problem (which may not be 1-to-1) to find the effective classifier. We discuss cases where this would and wouldn’t work in Sec. 4.4.
  2. 2.
    Regularizing curvature: A sufficient condition for a pre-image strategic $h^{\prime}$ to exist for a given non-strategic $h$ is that the function mapping is 1-to-1. Rather than restrict learning to only invertible classes $H^{\prime}$ a priori, an alternative approach is to work with general $H^{\prime}$ but regularize against those $h^{\prime}\in H^{\prime}$ that do not have an inverse. Prop. 1 implies that one way to promote this is by encouraging $h^{\prime}$ that are smooth, since low-curvature classifiers are less prone to direct wipeout, and therefore more likely to permit inversion.
  3. 3.
    Finding the interpolation threshold: The interpolation threshold — the point in which the number of model parameters (or model complexity more generally) attains minimal training error — is key for learning in practice: it marks the extreme point of overfitting (in the classic underparametrized regime) and the beginning of benign overfitting (in the modern overparametrized regime, e.g., of deep neural nets). In non-strategic learning, this point is easily identified as that where the training error is zero. In contrast, our results show that in strategic learning, the minimal training error can be strictly positive. This makes it unclear when (and if) the threshold has been reached. Fortunately, our procedure for our approximation experiment in Sec. 6.2 can be used generally to compute the minimal attainable strategic training error, independently of the chosen model class. This provides a tool that can be used in practice to measure interpolation.
In terms of model class selection, our results in Sec. 5 offer insight into potential upsides and downsides to certain classes. One implication of our results is that strategic behavior can generally either increase or decrease the complexity of the chosen model class. However, Thm. 2 ensures that the VC dimension can only increase for any class that is closed to input scaling, which includes many common classes used in practice. Thus, a learner who chooses such classes is guaranteed that the potential expressivity of the learned classifier will not be deteriorated by strategic behavior. Additionally, Prop. 4, Thm. 4, and Thm. 3 all give upper bounds on the SVC of particular classes (e.g., ReLU neural networks). The implication is that a learner choosing to work with such $H$ is guaranteed that if $H$ is learnable, then the induced $H_{\Delta}$ is also learnable. Finally, Obs. 5 and Thm. 1 both point to the fact that several non-strategic classifiers $h$ can be mapped to the exact same effective classifier $h_{\Delta}$. If a class $H$ includes many such cases, then it has significant redundancy, which can have negative implications on the process of choosing a good $h\in H$ via learning (e.g., in terms of generalization, optimization, or the effectiveness of proxy losses). Ideally, the learner should be cautious of classes in which this is prevalent, for example, classes that permit high curvature or small negative areas.
@@ D.1 Experiment #1: Expressivity
In Sec. 6.1, we experimentally demonstrate how the strategic setting affects classifier expressivity. In this experiment, we first randomly sample a degree-$k$ polynomial $h$, then determine the set of points $S$ that make up the decision boundary of $h_{\Delta}$ (i.e., are directly adjacent to it), and finally find the best approximate polynomial fit of $S$  [^8] Because we are concerned specifically with the polynomial fit of $h_{\Delta}$, we label the points in a fine grid G both regularly and then strategically, and use the strategic labels to get a polynomial fit. In this way, we do not directly compute the effective classifier (which would be difficult), but approximate it well enough to get the polynomial fit as well as visualize it.. Because user features are often naturally bounded, we assume that $x\in{\cal{X}}=\mathbb{R}^{2}$ and that all $x$ are bounded by $|x|_{\infty}\leq 100$ (see Appendix A.2 for a discussion on how our theoretical results hold under bounded input domains). We report average values and standard error confidence intervals of 100 instances of the setup for each $\alpha\in\{2,4,8,16,24,32,40\}$ and $k\in[1,10]$. The instances were divided among 100 CPUs to speed up computation.
To randomly sample polynomials of degree $k$, we leverage the fact that the number of points needed to uniquely determine a degree-$k$ polynomial over $\mathbb{R}^{2}$ is $k+2\choose 2$. Therefore, for each instance, we randomly choose $k+2\choose 2$ points and labels, and use an SVM polynomial fitter to find the best fitting degree-$k$ polynomial fit to these points. To ensure that the generated polynomial is indeed a full-degree polynomial, we construct a grid $G$ of points over $100,100]^{2}$, label each point $g\in G$ using $h$ to get labels $Y$, and verify that the SVM best-fitting degree-$k$ polynomial achieves near-perfect accuracy on ($G,Y$), while the best (k1)(k-1)-degree polynomial achieves low accuracy on ($G,Y$).
@@ D.2 Experiment #2: Approximation
In Sec. 6.2, we experimentally demonstrate the effects of non-universality on the maximum strategic accuracy. This experiment consists of (i) sampling synthetic data and (ii) calculating the maximum linear accuracy and strategic accuracy on each dataset instance. We report average values and standard error confidence intervals of 20 instances of the setup for each $\alpha\in\{0.5,1,2\}$ and $\mu\in[0,5]$. The instances were divided among 100 CPUs to speed up computation.
@@
To calculate the maximum linear accuracy in each instance, we first note that there exists an optimal linear classifier that runs through (or infinitesimally close to for negative points) exactly two dataset points. This is because for any optimal classifier $h$ that runs through exactly one dataset point $x\in{\cal{X}}$, $h$ may be rotated until it runs through another $x^{\prime}\in{\cal{X}}$ without changing any of the labels. For any optimal classifier $h$ that runs through no $x\in{\cal{X}}$, $h$ can be shifted until it runs through one (or two collinear) $x\in{\cal{X}}$ and then rotated appropriately. Because all data points are random, we assume that no three points are collinear. Therefore, the maximum linear accuracy on $\cal{X}$} can be calculated by taking the maximum accuracy over the $n\choose 2$ classifiers that run through each pair of $x,x^{\prime}\in{\cal{X}}$  [^9] Because points on the decision boundary are labeled as positive, we manually define that negative points used to define the linear classifier are accurately labeled to reflect the fact that the optimal classifier does not run through the point, but infinitely close to it..
When calculating the maximum strategic accuracy, we note that the minimum size of a positive region in $h_{\Delta}$ is the ball $B_{\alpha}$, so any strategic classifier can be represented by the intersection of multiple $B_{\alpha}$. We therefore calculate the set $S$ of all possible subsets $s\subseteq{\cal{X}}$ that any $B_{\alpha}$ classifies as positive, and find the intersection $\bigcap_{s\in S^{\prime}\subseteq S}\,s$ with the highest accuracy. To find all $s\in S$, we iterate through the circles centered at each point on a fine-grained grid over the range of the dataset (padded by $\alpha$ on all sides) and check which points $x\in{\cal{X}}$ each circle includes. To find the best intersection of these circles, we iterate through each of the subsets of $Z\subseteq{\cal{X}}_{0}$  [^10] Subsets are traversed in size order with early stopping once the size of the subset is large enough that the accuracy will be worse than the current best accuracy, even if all positive points are correctly classified. and then $s\in S$ to find the maximum number of positive points that can be classified as positive if we allow ourselves to mistakenly classify the negative points in $Z$.